OpenStudy (anonymous):

how do you factor x^2-4x+24

OpenStudy (anonymous):

Find the roots :-D \[x^2 - 4x + 4 = -20\] \[(x-2)^2 = -20 \]The above statement is false, so you cannot factor x^2-4x+24

OpenStudy (anonymous):

If you must factor it, however, you may use complex numbers.

OpenStudy (anonymous):

its not factorable, u can try using the quadratic formula

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\[x - 2=i\sqrt{20}=2\sqrt{5}*i\]or\[x-2=-i\sqrt{20}=-2\sqrt{5}*i\]

OpenStudy (anonymous):

It may therefor be factored into\[(x - (2-2\sqrt{5}*i))(x-(2+2\sqrt{5}*i))\]